(y-zx)p+(x+yz)q=x2+y2

1
Expert's answer
2020-08-03T18:48:07-0400

"\frac{dx}{y-xz}=\frac{dy}{x+yz}=\frac{dz}{x^2+y^2}=t\\
dx=yt-xzt,\quad dy=xt+yzt,\quad dz=(x^2+y^2)t\\
ydx+xdy=y^2t-xyzt+x^2t+xyzt=(x^2+y^2)t\Longrightarrow\\
\frac{ydx+xdy}{x^2+y^2}=\frac{dz}{x^2+y^2}=t\Longrightarrow\\
dz=ydx+xdy\Longrightarrow z'_x=y,\quad z'_y=x \Longrightarrow z(x,y)=xy+c;\\
check:\\
(y-xz)p+(x+yz)q=(y-x\cdot (xy+c))\cdot z'_x + (x+y\cdot (xy+c))\cdot z'_y=\\
y^2-xy(xy+c)+x^2+xy(xy+c)=x^2+y^2;"